OpenStudy (anonymous):

Which of the following is the simplified form of fifth root of x times the fifth root of x times the fifth root of x times the fifth root of x?

OpenStudy (anonymous):

\[\sqrt[5]{x}^4=x^{4/5} \]

OpenStudy (anonymous):

so you leave the exponent as the same number no adding or subtracting

OpenStudy (anonymous):

\[\sqrt[5]{x} \sqrt[5]{x} \sqrt[5]{x} \sqrt[5]{x}=x^{4/5} \]